# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ Abstract.
Recently, foundation language models (LMs) have marked significant achievements in the domains of natural language processing (NLP) and computer vision (CV). Unlike traditional neural network models, foundation LMs obtain a great ability for transfer learning by acquiring rich commonsense knowledge through pre-training on extensive unsupervised datasets with a vast number of parameters. However, they still can not emulate human-like continuous learning due to catastrophic forgetting. Consequently, various continual learning (CL)-based methodologies have been developed to refine LMs, enabling them to adapt to new tasks without forgetting previous knowledge. However, a systematic taxonomy of existing approaches and a comparison of their performance are still lacking, which is the gap that our survey aims to fill. We delve into a comprehensive review, summarization, and classification of the existing literature on CL-based approaches applied to foundation language models, such as pre-trained language models (PLMs), large language models (LLMs) and vision-language models (VLMs). We divide these studies into offline CL and online CL, which consist of traditional methods, parameter-efficient-based methods, instruction tuning-based methods and continual pre-training methods. Offline CL encompasses domain-incremental learning, task-incremental learning, and class-incremental learning, while online CL is subdivided into hard task boundary and blurry task boundary settings. Additionally, we outline the typical datasets and metrics employed in CL research and provide a detailed analysis of the challenges and future work for LMs-based continual learning.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Recent advancements in foundation language models (LMs) have set new benchmarks in both natural language processing (NLP) (Zhao et al., 2023; Min et al., 2023; Zhou et al., 2023a) and computer vision (CV) (Wang et al., 2023b). Foundation LMs encompass three primary categories: Pre-trained Language Models (PLMs) (Min et al., 2023), Large Language Models (LLMs) (Zhao et al., 2023), and Vision-Language Models (VLMs) (Du et al., 2022). PLMs such as BERT (Kenton and Toutanova, 2019), RoBERTa (Liu et al., 2019a), and BART (Lewis et al., 2020) focus on text-based tasks and are crucial for understanding and generating language by leveraging tasks like masked language modeling during pre-training. LLMs, including models like GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), extend the capabilities of PLMs by increasing the scale of model architecture and training data, thus enhancing their generality and adaptability across a broader range of tasks. VLMs, represented by VisualBERT (Li et al., 2019), CLIP (Radford et al., 2021), LLaVA (Liu et al., 2024) and DALL-E (Ramesh et al., 2021), integrate text and image modalities to enable complicated interactions between visual and textual information. The underlying paradigm of these models involves pre-training on extensive, often unlabeled datasets to capture rich semantic information, which is subsequently fine-tuned for specific tasks or domains. This methodology not only boosts performance across various applications but also significantly enhances the models’ flexibility and task adaptability.
@@
However, these foundation models often demonstrate limitations in dynamic environments with a sequence of tasks, primarily due to their fixed parameters once training is completed. These models generally lack the capability to integrate new data or concepts without undergoing a retraining process. A significant challenge associated with training on a sequence of tasks is “catastrophic forgetting” (Kirkpatrick et al., 2017), a phenomenon where a model loses previously acquired knowledge upon learning new information. This is in stark contrast to human learning processes, which are inherently continuous and adaptive. Despite the successes of multi-task learning (MTL) and transfer learning (TL) in certain applications, they have limitations in real-world scenarios. MTL necessitates having all tasks and their data available upfront, which poses a challenge when launching a new service as the model must be retrained with all the data. Furthermore, TL is typically done with only two tasks, i.e., the source and the target, rendering it impractical for real-world online platforms with multiple target tasks. To address these challenges, it is crucial for models to process and learn the continuously expanding and diversifying datasets. This requires mechanisms that allow models to adapt to new linguistic phenomena and trends without compromising the accuracy and sensitivity towards historical data.
Consequently, continual learning (CL) (Wang et al., 2024b; Van de Ven et al., 2022), also referred to as lifelong learning (Parisi et al., 2019) or incremental learning (Zhou et al., 2023b), is a crucial area in artificial intelligence that seeks to develop systems capable of continuously updating themselves and acquiring new knowledge, without forgetting previously learned information, similar to human learning (De Lange et al., 2021). This paradigm is particularly relevant in the context of foundation language models (LMs), which are challenged by specific issues such as catastrophic forgetting (CF) and cross-task knowledge transfer (KT). Catastrophic forgetting represents a significant challenge, where a model tends to lose previously acquired knowledge upon learning new information. To address this, language models must maintain a robust grasp of past language data while adapting to new linguistic trends. Furthermore, cross-task knowledge transfer is essential for enhancing the continual learning process. Effective KT not only accelerates the learning curve for new tasks (forward transfer) but also enhances the model’s performance on prior tasks via the feedback of new knowledge (backward transfer).
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=10.0em,font=,, where level=4text width=10.5em,font=,, [ Foundation LMs-based CL, ver [ Offline Continual
@@
CPT (Ke et al., 2022), C-PT (Zhu et al., 2022), CL-KD (Castellucci et al., 2021), PlugLM (Cheng et al., 2023b), Pretr (Cossu et al., 2022), AEWC (Lee, 2017), Continual DAP-training (Ke et al., 2023) , leaf, text width=44.6em ] ] [ LLM-based DIL  (§4.1.2) [ COPR (Zhang et al., 2023b), LAMOL (Sun et al., 2020a), RVAE_LAMOL (Wang et al., 2022a), Adapt-Retrieve-Revise (Zhang et al., 2023c), Lifelong-MoE (Chen et al., 2023),
CKL (Jang et al., 2022b), DACP (Xie et al., 2023), CPPO (Zhang et al., 2024), EcomGPT-CT (Ma et al., 2023) , leaf, text width=44.6em ] ] [ VLMs-based DIL  (§4.1.3) [ S-Prompt (Wang et al., 2022b), Medical AI (Yi et al., 2023), VQACL (Zhang et al., 2023d) , leaf, text width=44.6em ] ] ] [ Task-Incremental
Learning  (§4.2) [ PLMs-based TIL  (§4.2.1) [ PP (Razdaibiedina et al., 2023), CTR (Ke et al., 2021a), MeLL (Wang et al., 2021a), LINC (Liu and Mazumder, 2021), ERDA (Qin and Joty, 2022a), PCLL (Zhao et al., 2022), CLIF (Jin et al., 2021a), ConTinTin (Yin et al., 2022),
HMI (Maekawa et al., 2023), Adaptive Compositional Modules (Zhang et al., 2022b), DYNAINST (Mok et al., 2023), Conure (Yuan et al., 2021), TERACON (Kim et al., 2023),
ERNIE 2.0 (Sun et al., 2020b), RecyclableTuning (Qin et al., 2023) , leaf, text width=44.6em ] ] [ LLMs-based TIL  (§4.2.2) [ Conpet (Song et al., 2023), InstructAlign (Cahyawijaya et al., 2023), Continual-T0 (Scialom et al., 2022a), DynaMind (Du et al., 2023), ELM (Jang et al., 2023), O-LoRA (Wang et al., 2023a), JARe (PENG et al., 2024) , leaf, text width=44.6em ] ] [ VLMs-based TIL  (§4.2.3) [ Medical AI (Yi et al., 2023), CTP (Zhu et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapters4CL (Yu et al., 2024), TRIPLET (Qian et al., 2023) , leaf, text width=44.6em ] ] ] [ Class-Incremental
@@
SRC (Liu et al., 2019b) , leaf, text width=44.6em ] ] [ VLMs-based CIL  (§4.3.2) [ MoE-Adapters4CL (Yu et al., 2024), VLM-PL (Kim et al., 2024), Adaptation-CLIP (Liu et al., 2023a), PROOF (Zhou et al., 2023c), LGCL (Khan et al., 2023), ZSCL (Zheng et al., 2023),
CLAP (Jha et al., 2024), GMM (Cao et al., 2024) , leaf, text width=44.6em ] ] ] ] [ Online Continual
@@
Boundary  (§5.1) [ PLMs-based HTB  (§5.1.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), ProgModel (Shen et al., 2019) , leaf, text width=44.6em ] ] ] [ Blurry Task
Boundary  (§5.2) [ PLMs-based BTB  (§5.2.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b) , leaf, text width=44.6em ] ] [ VLMs-based BTB  (§5.2.2) [ CBA (Wang et al., 2023d), MVP (Moon et al., 2023), DKR (Cui et al., 2024) , leaf, text width=44.6em ] ] ] ] ]
@@
Recent advancements in continual learning methodologies have substantially enhanced the adaptability and knowledge retention capabilities of foundational language models (LMs). These developments are crucial for addressing complex challenges previously observed in CL. Researchers have formulated innovative strategies to mitigate these challenges, thereby enabling LMs to maintain high performance across a variety of tasks while continually integrating new knowledge (Mehta et al., 2021; Cossu et al., 2022; Lee et al., 2023). Notable successes have been documented in diverse downstream tasks, such as aspect-based sentiment analysis, where continual learning enables dynamic adaptation to evolving aspects and sentiments (Ke et al., 2021b). Similarly, in dialogue generation, the novel technologies assist models in refining and expanding their conversational capabilities through ongoing interactions (Scialom et al., 2022b). In text classification, continual learning facilitates the incorporation of new categories and adjustments to shifts in text distributions without the need for complete retraining (Razdaibiedina et al., 2023). Moreover, in the realm of visual question answering, continual learning is essential for updating the models’ capabilities to process and respond to new types of visual content and queries (Zhang et al., 2023d; Qian et al., 2023). The aforementioned works underscore the potential of continual learning to significantly boost the performance of foundation LMs.
@@
In the domain of continual learning, there has been a significant paradigm shift from traditional methodologies to those that integrate foundation LMs (See Figure 1). First, foundation LMs demonstrate enhanced generalization and transfer learning abilities across diverse tasks owing to their broad pre-training on large-scale datasets. The model has specialized transfer capability to quickly adapt to downstream tasks with only a few samples. Consequently, it is crucial to mitigate the degradation of both the zero-shot transfer and history task abilities in LMs while facilitating the acquisition of new skills. Second, due to the substantial number of parameters in foundation LMs, it is crucial to employ parameter-efficient techniques (Han et al., 2024), such as prompt tuning (Liu et al., 2022) and adapters (Mundra et al., 2024), to update parameters without comprehensive retraining. Third, the foundation LMs possess the capability to follow instructions through instructional learning (Dong et al., 2022; Ouyang et al., 2022), enabling more dynamic and context-aware interactions.
@@
This review systematically categorizes these strategies and technologies into two core areas: offline continual learning and online continual learning (Figure 2). We first give detailed definitions and scenarios to format the setting of offline and online CL, where offline CL includes domain-incremental, task-incremental and class-incremental CL, and online CL includes hard task boundary and blurry task boundary. These learning strategies are further subdivided into methods based on Pre-trained Language Models (PLMs), Large Language Models (LLMs), and Vision-Language Models (VLMs). Then, we summarize the related papers about traditional methods, continual pre-training methods, parameter-efficient tuning methods and instruction-based methods. Finally, we static the main datasets from various perspectives and review the key metrics to evaluate the forgetting and transferring of the models.
The main contributions of this survey paper can be summarized as follows.
@@
  • We provide the characters of existing datasets for CL and present the main metrics to evaluate the performance of preventing forgetting and knowledge transfer.
  • We discuss the most challenging problems of foundation LMs-based CL and point out promising future research directions in this field.
The paper is organized as follows. In Section 2, we review the mainly related surveys about continual learning. Then, we introduce the base settings and learning modes of continual learning in Section 3, including the definitions and scenarios of CL. Furthermore, we present the related studies about offline continual learning, which can be divided into domain-incremental learning, task-incremental learning and class-incremental learning in Section 4. In Section 5, we focus on online continual learning, including hard task boundary and blurry task boundary settings. The typical datasets and metrics are provided in Section 6 and 7. Finally, we analyze the challenge and further work in Section 8 and give the conclusion in Section 9.
@@ 2.1. Continual Learning
Early examinations of Continual Learning (CL) have provided broad coverage, as observed in surveys such as Parisi et al. (Parisi et al., 2019). Recently, Wang et al. (Wang et al., 2024b) conduct a comprehensive survey that categorizes five key strategies in CL: regularization-based, replay-based, optimization-based, representation-based, and architecture-based approaches. This survey reflects an effort to organize and understand the diverse methodologies employed in the field. Notably, there is a growing focus on class-incremental setting (Zhou et al., 2023b; Masana et al., 2022; Belouadah et al., 2021) and replay-based approaches (Hayes et al., 2021), reflecting the increasing granularity of research interests within the CL domain.

### v2

@@ Abstract.
Recently, foundation language models (LMs) have marked significant achievements in the domains of natural language processing (NLP) and computer vision (CV). Unlike traditional neural network models, foundation LMs obtain a great ability for transfer learning by acquiring rich commonsense knowledge through pre-training on extensive unsupervised datasets with a vast number of parameters. Despite these capabilities, LMs still struggle with catastrophic forgetting, hindering their ability to learn continuously like humans. To address this, continual learning (CL) methodologies have been introduced, allowing LMs to adapt to new tasks while retaining learned knowledge. However, a systematic taxonomy of existing approaches and a comparison of their performance are still lacking. In this paper, we delve into a comprehensive review, summarization, and classification of the existing literature on CL-based approaches applied to foundation language models, such as pre-trained language models (PLMs), large language models (LLMs) and vision-language models (VLMs). We divide these studies into offline and online CL, which consist of traditional methods, parameter-efficient-based methods, instruction tuning-based methods and continual pre-training methods. Additionally, we outline the typical datasets and metrics employed in CL research and provide a detailed analysis of the challenges and future work for LMs-based continual learning.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Recent advancements in foundation language models (LMs) have set new benchmarks in both natural language processing (NLP) (Zhao et al., 2023; Min et al., 2023; Zhou et al., 2023a) and computer vision (CV) (Wang et al., 2023b; Tiwari et al., 2022; Ramasesh et al., 2022). Foundation LMs encompass three primary categories: Pre-trained Language Models (PLMs) (Min et al., 2023), Large Language Models (LLMs) (Zhao et al., 2023), and Vision-Language Models (VLMs) (Du et al., 2022). These models are pre-trained on large, unlabeled datasets to capture rich semantic information, which is then fine-tuned for specific tasks or domains. This strategy not only enhances performance across various applications but also significantly improves the flexibility and adaptability of these models.
However, despite their strengths, foundation LMs face challenges in dynamic environments where tasks evolve over time. A key issue is “catastrophic forgetting” (Kirkpatrick et al., 2017), where models lose previously learned knowledge when adapting to new information. Unlike human learning, which is inherently continuous and adaptive (De Lange et al., 2021), foundation LMs generally require retraining to incorporate new data. Effective learning in such environments demands not only the ability to accelerate learning on new tasks (forward transfer) but also to improve performance on previous tasks by integrating newly acquired knowledge (backward transfer). While multi-task learning (MTL) and transfer learning (TL) offer potential solutions, MTL requires all task data to be available upfront, and TL focuses on limited tasks, making both approaches impractical for dynamic, real-world applications.
@@
Continual learning (CL) (Wang et al., 2024c; Van de Ven et al., 2022), also known as lifelong learning (Parisi et al., 2019) or incremental learning (Zhou et al., 2023b), offers an effective solution to these challenges. It aims to develop systems capable of continuously learning and updating without forgetting past knowledge. Recent advancements in CL methodologies have substantially enhanced the adaptability and knowledge retention capabilities of foundation LMs (Mehta et al., 2021; Cossu et al., 2022; Lee et al., 2023). Notable successes have been documented in diverse downstream tasks, such as aspect-based sentiment analysis (Ke et al., 2021b), dialogue generation (Scialom et al., 2022a), text classification (Razdaibiedina et al., 2023), visual question answering (Zhang et al., 2023d; Qian et al., 2023) and so on. Luo et al. (Luo et al., 2023) conduct an empirical study on catastrophic forgetting (CF) in large language models (LLMs) during continual instruction tuning. The aforementioned works underscore the potential of continual learning to significantly boost the performance of foundation LMs.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=10.5em,font=,, where level=4text width=10.5em,font=,, [ Foundation LMs-based CL, ver [ Offline Continual
@@
Pretr (Cossu et al., 2022), CPT (Ke et al., 2022), C-PT (Zhu et al., 2022), CL-KD (Castellucci et al., 2021), PlugLM (Cheng et al., 2023b), AEWC (Lee, 2017), DAS (Ke et al., 2023) , leaf, text width=44.6em ] ] [ LLM-based DIL  (§5.1.2) [ COPF (Zhang et al., 2023b), LAMOL (Sun et al., 2020a), RVAE_LAMOL (Wang et al., 2022a), Adapt-Retrieve-Revise (Zhang et al., 2023c), Lifelong-MoE (Chen et al., 2023),
DACP (Xie et al., 2023), CPPO (Zhang et al., 2024b), EcomGPT-CT (Ma et al., 2023), LLM-CL (Ding et al., 2024), AMA (Lin et al., 2024) , leaf, text width=44.6em ] ] [ VLMs-based DIL  (§5.1.3) [ S-Prompt (Wang et al., 2022b), VQACL (Zhang et al., 2023d), SC-MLLM (Liu et al., 2024), DIKI (Tang et al., 2024) , leaf, text width=44.6em ] ] ] [ Task-Incremental
Learning  (§5.2) [ PLMs-based TIL  (§5.2.1) [ PP (Razdaibiedina et al., 2023), CTR (Ke et al., 2021a), MeLL (Wang et al., 2021), LINC (Liu and Mazumder, 2021), ERDA (Qin and Joty, 2022a), PCLL (Zhao et al., 2022), BiHNet-Reg (Jin et al., 2021), ConTinTin (Yin et al., 2022),
HMI (Maekawa et al., 2023), ACM (Zhang et al., 2022), DYNAINST (Mok et al., 2023), Conure (Yuan et al., 2021), TERACON (Kim et al., 2023),
EMR (Wang et al., 2019b), ERNIE 2.0 (Sun et al., 2020b), RecyclableTuning (Qin et al., 2023), RecAdam (Chen et al., 2020) , leaf, text width=44.6em ] ] [ LLMs-based TIL  (§5.2.2) [ Conpet (Song et al., 2023), InstructAlign (Cahyawijaya et al., 2023), Continual-T0 (Scialom et al., 2022b), DynaMind (Du et al., 2023), ELM (Jang et al., 2023), O-LoRA (Wang et al., 2023a), JARe (PENG et al., 2024),
Robocoder (Li et al., 2024), Eureka (Ma et al., 2024), COPF (Zhang et al., 2024a) , leaf, text width=44.6em ] ] [ VLMs-based TIL  (§5.2.3) [ Medical AI (Yi et al., 2023), CTP (Zhu et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapters4CL (Yu et al., 2024b), TRIPLET (Qian et al., 2023), AwoForget (Zheng et al., [n. d.]),
SND (Yu et al., 2024a) , leaf, text width=44.6em ] ] ] [ Class-Incremental
@@
SRC (Liu et al., 2019b) , leaf, text width=44.6em ] ] [ VLMs-based CIL  (§5.3.2) [ MoE-Adapters4CL (Yu et al., 2024b), VLM-PL (Kim et al., 2024), Adaptation-CLIP (Liu et al., 2023a), PROOF (Zhou et al., 2023c), LGCL (Khan et al., 2023), ZSCL (Zheng et al., 2023),
CLAP (Jha et al., 2024), GMM (Cao et al., 2024), RAPF (Huang et al., 2024), STAR-Prompt (Menabue et al., 2024), DIKI (Tang et al., 2024), SND (Yu et al., 2024a), AwoForget (Zheng et al., [n. d.]) , leaf, text width=44.6em ] ] ] ] [ Online Continual
@@
Boundary  (§6.1) [ PLMs-based HTB  (§6.1.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), ProgModel (Shen et al., 2019) , leaf, text width=44.6em ] ] [ VLMs-based HTB  (§6.1.2) [ PEGP (Qiao et al., 2024) , leaf, text width=44.6em ] ] ] [ Blurry Task
Boundary  (§6.2) [ PLMs-based BTB  (§6.2.1) [ MBPA++ (de Masson D’Autume et al., 2019), Meta-MBPA++ (Wang et al., 2020), OML-ER (Holla et al., 2020), TPEM (Geng et al., 2021), CID (Liu et al., 2021b), S6 (Li et al., 2023) , leaf, text width=44.6em ] ] [ VLMs-based BTB  (§6.2.2) [ DKR (Cui et al., 2024), SIT (Wang et al., 2024a), OLiVia-Nav (Narasimhan et al., 2024), G-NoCL (Seo et al., 2024) , leaf, text width=44.6em ] ] ] ] ]
@@
In the domain of continual learning, there has been a shift from traditional methods to those incorporating foundation LMs (Figure 1). First, foundation LMs has specialized transfer capability to quickly adapt to downstream tasks with only a few samples. Consequently, it is crucial to mitigate the degradation of both the zero-shot transfer and history task abilities while facilitating the acquisition of new skills. Second, due to the substantial number of parameters in foundation LMs, it is crucial to employ parameter-efficient techniques (Han et al., 2024), such as prompt tuning (Liu et al., 2022) and adapters (Mundra et al., 2024), to update parameters without comprehensive retraining. Third, the foundation LMs possess the capability to follow instructions through instructional learning (Dong et al., 2022; Ouyang et al., 2022), enabling more dynamic and context-aware interactions.
@@
This review systematically organizes continual learning strategies and technologies into two main categories: offline continual learning and online continual learning (Figure 2). We begin by defining and explaining the different settings for these two types of continual learning. Offline continual learning includes domain/task/class-incremental CL, while online continual learning is further divided into methods that address hard task boundaries and those that manage blurry task boundaries (Section 4). To further clarify the relationship between these strategies and model architectures, we group the methods based on three major model types: Pre-trained Language Models (PLMs), Large Language Models (LLMs), and Vision-Language Models (VLMs). These categories reflect the distinct requirements for continual learning in each architecture and emphasize the role these models play in shaping learning strategies. Additionally, we classify the methods into four key approaches: traditional continual learning methods, continual pre-training methods, parameter-efficient tuning methods, and instruction-based methods (Section 5, 6). This categorization highlights the different techniques used to address challenges like catastrophic forgetting and to improve knowledge transfer in continual learning scenarios. Finally, we static the main datasets from various perspectives (Section7) and review the key metrics to evaluate the forgetting and transferring of the models (Section 8).
The main contributions of this survey paper can be summarized as follows:
@@
  • We provide the characters of existing datasets for CL and present the main metrics to evaluate the performance of preventing forgetting and knowledge transfer. Furthermore, we discuss the most challenging problems of foundation LMs-based CL and point out promising future research directions in this field.
@@ 2.1. Continual Learning
Early examinations have provided broad coverage, as observed in surveys such as Parisi et al. (Parisi et al., 2019). Recently, Wang et al. (Wang et al., 2024c) conduct a comprehensive survey that categorizes five key strategies in CL: regularization-based, replay-based, optimization-based, representation-based, and architecture-based approaches. This survey reflects an effort to organize and understand the diverse methodologies employed in the field. Notably, there is also a growing focus on class-incremental setting (Zhou et al., 2023b; Masana et al., 2022; Belouadah et al., 2021) and replay-based approaches (Hayes et al., 2021). Recent surveys (Wu et al., 2024; Shi et al., 2024; Zheng et al., 2024) provide comprehensive overviews of continual pre-training, continual domain-adaptive pre-training, continual instruction tuning, and continual alignment. In particular, Shi et al. (Shi et al., 2024) provides an in-depth overview of CL, specifically in LLMs, introducing vertical CL (adapting from general to specific domains) and horizontal CL (adapting across time and domains). Additionally, one notable approach in continual alignment is Reinforcement Learning from Human Feedback (RLHF) (Wu et al., 2022), which incorporates human feedback to optimize model outputs for better task adaptation.